OpenStudy (anonymous):

Im graphing a parabola using the function; y = 3x^2 - 6x + 1. My question says what is the equation of the axis of symmetry, how would I solve it? (simple easy version plzz) xD

OpenStudy (anonymous):

make a sketch of the function from -3 to +3 in intervals of 1. Then think about the questions again it should become more clear

OpenStudy (amistre64):

-b/2a

OpenStudy (amistre64):

-middle#/2(first#) include signs

OpenStudy (amistre64):

x = --6/2(3) = 6/6 = 1

OpenStudy (anonymous):

oh sorry i did not reply hold on let me read and try to comprehend what you wrote.

OpenStudy (anonymous):

now looking at this you could also solve it graphically if that helps understand why amistre's equation is what it is

OpenStudy (anonymous):

how would you solve it grapically then, i am having a hard time putting the info together.... ;P

OpenStudy (anonymous):

ok so when they say axis of symmetry the mean, graphically, where on the graph could you make a line so that if you put a mirror on that line you would see a mirror image of whatever is on the other side. so here it was y=1, which is a straight line going up and down from x=1. so if you put a mirror on that line, looking at the graph, would you see the mirror image?

OpenStudy (anonymous):

I entered the equation into the graphing calculator and ok yea its y = 1 right.... Yes you would wouldnt you?...

OpenStudy (anonymous):

see a mirror image. that is

OpenStudy (anonymous):

i dunno that is something you cant explain, you just have to look at the graph and see what i mean since it is really obvious in this case. just think about it and you will get it :)

OpenStudy (anonymous):

haha blonde here!! But ya... maybe in a few days i'll figure it out jk lol (:

OpenStudy (anonymous):

ok ya just looked at it and im pretty sure i figured it out... thx!! (: haha not a math kinda girl im a writer.. :P but anyways thx!!